# Similarity of Neural Architectures using Adversarial Attack Transferability

### v3

@@ 4 Model analysis by network similarity
We quantitatively analyze how each model component affects the similarity or diversity between networks by feature importance analysis and clustering analysis. We also show that the diversity caused by different training strategies is not as significant as caused by different neural architectures.
4.1 Which architectural component causes the difference between architectures?
Figure 3: Importance of architectural components to network similarity. 13 components are sorted by the contribution to the network similarities. The larger feature importance means the component contributes more to the network similarity.
@@ 4.1 Which architectural component causes the difference between architectures?
Feature important analysis. The full list of components of 69 architectures is in Appendix. Now, we measure the feature importance by fitting a gradient boosting regressor [30] on the feature difference (e.g., $f_{\text{ResNet-50}}-f_{\text{DeiT-base}}$) measured by Hamming distance and the corresponding similarity. The details of the regressor are described in Appendix. We use the permutation importance [7] that indicates how the trained regression model changes the prediction according to randomly changing each feature. The feature importance of each architectural component is shown in Fig. 3. We first observe that the choice of base architecture (e.g., CNN [57], Transformer [101], and MLP-Mixer [95]) contributes to the similarity most significantly. Fig. 3 also shows that the design choice of the input layer (i.e., stem layer design choice or input resolution) affects the similarity as much as the choice of basic components such as normalization layers, activation functions, and the existence of attention layers. On the other hand, we observe that the modified efficiency-aware convolution operations, such as depth-wise convolution [17] or group convolution [110], are ineffective for model diversity.
@@ 4 Stem layer: stack of 3$\times$3 conv, 2D SA, HaloNet-50 [102], LambdaResNet-50 [5], BoTNeT-26 [87],
6 Stem layer: 7$\times$7 conv w/ s2, ReLU, ResNet-50 [40], ResNet-101 [40], ResNeXt-50 [110], Wide ResNet-50 [116],
Pooling at stem, CNN, BN SE-ResNet-50 [46], SE-ResNeXt-50 [46], ResNet-V2-50 [41], ResNet-V2-101 [41],
ResNet-50 (GN) [109], ResNet-50 (BlurPool) [120], DPN-107 [14], Xception-65 [17]
7 NAS, Stem layer: 3$\times$3 conv w/ s2 EfficientNet-B2 [93], FBNetV3-G [23], ReXNet ($\times$1.5) [38], RegNetY-32 [76],
CWA: middle of blocks, CWA, DW Conv MixNet-XL [94], NF-RegNet-B1 [8]
@@
8 Input resolution: 224, Stem layer: stack of 3$\times$3 convs, NFNet-L0222Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63]., ECA-NFNet-L0222Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63]., PoolFormer-M48 [113], ResNeSt-50 [118], ResNet-V2-50-D-EVOS [^2] Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63].,
Group Conv, Final GAP, 2D SA ConvMixer-1536/20 [99]
9 ReLU, Input resolution: 224, DW Conv, BN, ViT-B (patch size 32) [28], R26+ViT-S [88], DLA-X-102 [112], eSE-VoVNet-39 [60],
2D self-attention ResNet-101-C [43], RegNetX-320 [76], HRNet-W32 [105]
10 ReLU + Leaky ReLU, InRes: 256, CSPResNet-50 [104], CSPResNeXt-50 [104], CSPDarkNet-53 [6], NF-ResNet-50 [8]
Stem layer: 7$\times$7 conv, CNN, Pooling at stem
Table 2: Clustering by similarities. We perform a clustering analysis on 69 architectures. All the architectures here are denoted by the aliases defined in their respective papers. We show the top-5 keywords for each cluster based on TF-IDF. InRes, SA, and CWA denote input resolution, self-attention, and channel-wise attention, respectively. The customized model details are described in the ${}^{\dagger}$.
@@
Clustering analysis. We additionally provide a clustering analysis based on the architectural similarities. We construct a pairwise similarity graph with adjacency matrix $A$ between all 69 architectures where its vertex denotes an architecture, and its edge denotes the similarity between two networks. We perform the spectral clustering [71] on $A$ where the number of clusters $K$ is set to 10: We compute the Laplacian matrix of $A$, $L=D-A$ where $D$ is the diagonal matrix and its $i$-th component is $\sum_{j}A_{ij}$. Then, we perform K-means clustering on the $K$-largest eigenvectors of $L$, where $K$ is the number of clusters. The pairwise distances of spectral features (i.e., 10-largest eigenvectors of $L$) of 69 neural architectures are shown in Fig. 3(a). The rows and columns of Fig. 3(a) are sorted by the clustering index (Tab. 2). We can see the block-diagonal patterns, i.e., in-clusters similarities are more significant than between-clusters similarities. More details are in Appendix.
@@
(a) Spectral feature
(b) Similarity score
Figure 4: Spectral feature from SAT of ResNet-50 and ViT-s with four different optimization settings. The rows and columns of (a) are sorted by the clustering index. More details with model names are described in Appendix. Init, LR, and WD are randomly chosen from models trained with different settings of initialization, learning rate, and weight decay in Tab. 3.
4.2 Training strategy and similarity score
The architectural difference is not the only cause of the model diversity. We compare the impact by different architecture choices (e.g., ResNet and ViT) and by different training strategies while fixing the model architecture, as follows: Different initializations can affect the model training by the nature of the stochasticity of the training procedure. For example, Somepalli et al. [85] showed that the decision boundary of each architecture could vary by different initializations. We also consider different optimization hyper-parameters (e.g., learning rate, weight decay). Finally, we study the effect of different training regimes (e.g., augmentations, type of supervision). For example, the choice of data augmentation (e.g., Mixup [117], or CutMix [114]) or label smoothing [92] can theoretically or empirically affect adversarial robustness [119, 72, 82, 19]. Similarly, we investigate the effect of different types of supervision, such as self-supervision (e.g., MoCo v3 [13], MAE [42] and BYOL [35]) or semi-weakly supervised learning [111]. Note that the training strategies mentioned later inevitably contain the earlier-mentioned training strategies. For example, when we train models with different training regimes, models have different initialization seeds and different optimization hyper-parameters. Comparing 69 different architectures also contains the effect of different initialization and optimization hyper-parameters and parts of different training regimes. This is necessary for achieving high classification performance. Experimental details about the used models are provided in Appendix.
Architecture Diversity by SAT
ResNet-50 Initialization 4.23
Hyper-parameter 4.05
Training regime 3.27
ViT-S Initialization 4.21
Hyper-parameter 4.22
Training regime 3.44
@@ 4.2 Training strategy and similarity score
Table 3: SAT within the same architecture. We compare the average similarity within the same architecture but trained with different procedures, e.g., random initializations, optimization hyper-parameters, and different training regimes. We report the similarities of two neural architectures, ResNet-50 and ViT-S. “All” denotes the average similarity of 69 neural architectures.
@@
Tab. 3 shows the comparison of similarity scores between the same architecture but different learning methods (a smaller similarity means more diversity). First, we observe that using different random initialization or different optimization hyper-parameters shows high correlations with each other (almost $\geq$ 4.2) while the average similarity score between various neural architectures is 2.73. In other words, the difference in initializations or optimization hyper-parameters does not significantly contribute to the model diversity. Second, we observe that using different learning techniques (e.g., different augmentation methods and types of supervision) remarkably affects the similarity score (3.27 for ResNet and 3.44 for ViT), but is not as significant as the architectural difference (2.73). Furthermore, the change of similarity caused by different initializations or hyper-parameters is less marked than the change caused by different architecture (Fig. 3(b)). These observations provide two insights into model similarity. First, they demonstrate that the diversity resulting from various training strategies is not significant enough to interfere with observing the diversity between various architectures. Second, constructing new architecture is a more efficient way to achieve varied models rather than re-training the same one. In the next section, we will show the need for diversity. The need for diversity is observed in the next section.
5 Analysis for Ensemble and Distillation
In this section, we analyze how our neural architecture similarity is related to downstream tasks involving more than one model. First, we show that using more diverse models will lead to better ensemble performance. Second, we study the relationship between knowledge distillation and SAT. Furthermore, we can suggest a similarity-based guideline for choosing a teacher model when distilling to a specific architecture. Through these observations, we can provide insights into the necessity of diverse models.
5.1 Model diversity and ensemble performance
@@
Figure 5: Correlation between SAT and ensemble performance. The trend line and its 90% confidence interval are shown.
@@
We first measure the 2-ensemble performances among the 69 neural architectures (i.e., the number of ensembles is $\binom{69}{2}=2346$). We plot the relationship between the model similarity and ERR in Fig. 5. We observe that there exists a strong negative correlation between the model similarity and the ensemble performance (Pearson correlation coefficient $-0.31$ with p-value $1.74\times 10^{-54}$), i.e., more diversity leads to better ensemble performance.
We also conduct $N$-ensemble experiments with $N\geq 2$ based on the clustering results (Tab. 2). We evaluate the average ERR of the ensemble of models from $k$ clusters, i.e., if $N=5$ and $k=3$, the ensembled models are only sampled from the selected 3 clusters while ignoring the other 7 clusters. We investigate the effect of model diversity and ensemble performance by examining $k=1\ldots N$ (i.e., larger $k$ denotes more diverse ensembled models). We report the result with the metric, error reduction rate (higher is better) in Fig. 5(a). ImageNet top-1 error of the ensemble is also shown in Appendix. In all metrics, we observe that the ensemble of more diverse models (i.e., by increasing the number of clusters for the model selection) shows better ensemble performance. Interestingly, in Fig. 5(a), we observe that when the number of clusters for the model selection ($k$) is decreased, the ensemble performance by the number of ensembled models ($N$) quickly reaches saturation. Similarly, in Fig. 5(b), we observe that the number of wrong samples by all models is decreased by selecting more diverse models. In other words, the different architectural choices can lead to different model decision boundaries.
(a) Error reduction rate
(b) “All wrong” samples ratio
Figure 6: Model diversity by architecture difference and ensemble performance. We report ensemble performances by varying the number of ensembled models ($N$) and the diversity of the ensembled models. The diversity is controlled by choosing the models from $k$ different clusters (larger $k$ denotes more diversity).
@@
Figure 7: Model diversity by training techniques and ensemble performance. The same metrics to Fig. 6 for ResNet-50 and ViT-S models in Tab. 3 are shown. “All” means $N$=$k$ in Fig. 6 (i.e., # of ensembled models = # of Clusters).
@@
Tab. 3 showed that the different training strategies are not as effective as different architectures in terms of diversity. To examine the effect of this observation on the ensemble scenario, we report the ensemble results of different training strategies, i.e., the same ResNet-50 and ViT-S models in Tab. 3. For comparison with different architectures, we also report the ensemble of different architectures where all ensembled models are from different clusters (i.e., when $N=k$ in Fig. 6). Fig. 7 shows that although using diverse training regimes (blue lines) improves ensemble performance most among other training techniques (red and green lines), the improvements by using different architectures (yellow lines) are more significant than the improvements by using different training regimes (blue lines) with large gaps.

### v4

@@ 4 Model Analysis by Network Similarity
4.1 Which Architectural Component Causes the Difference?
Settings.
@@ 4.1 Which Architectural Component Causes the Difference?
Feature important analysis.
Now, we measure the feature importance by fitting a gradient boosting regressor [32] on the feature difference (e.g., $f_{\text{ResNet-50}}-f_{\text{DeiT-base}}$) measured by Hamming distance and the corresponding similarity. The details of the regressor are described in Appendix. We use the permutation importance [9] that indicates how the trained regression model changes the prediction according to randomly changing each feature. The feature importance of each architectural component is shown in Fig. 4. We first observe that the choice of base architecture (e.g., CNN [60], Transformer [106], and MLP-Mixer [99]) contributes to the similarity most significantly. Fig. 4 also shows that the design choice of the input layer (i.e., stem layer design choice or input resolution) affects the similarity as much as the choice of basic components such as normalization layers, activation functions, and the existence of attention layers. On the other hand, we observe that the modified efficiency-aware convolution operations, such as depth-wise convolution [19], are ineffective for diversity.
Table 1: Clusters by SAT. All the architectures here are denoted by the aliases defined in their respective papers. We show the top-5 keywords for each cluster based on TF-IDF. InRes, SA, and CWA denote input resolution, self-attention, and channel-wise attention, respectively. The customized model details are described in the footnote.
@@ 4 Stem layer: stack of 3$\times$3 conv, 2D SA, HaloNet-50 [107], LambdaResNet-50 [7], BoTNeT-26 [92],
6 Stem layer: 7$\times$7 conv w/ s2, ReLU, ResNet-50 [42], ResNet-101 [42], ResNeXt-50 [117],
Pooling at stem, CNN, BN Wide ResNet-50 [124], SE-ResNet-50 [48], SE-ResNeXt-50 [48],
ResNet-V2-50 [43], ResNet-V2-101 [43], ResNet-50 (GN) [116],
ResNet-50 (BlurPool) [130], DPN-107 [16], Xception-65 [19]
7 NAS, Stem layer: 3$\times$3 conv w/ s2 EfficientNet-B2 [97], FBNetV3-G [24], ReXNet ($\times$1.5) [40],
CWA: middle of blocks, CWA, DW Conv RegNetY-32 [81], MixNet-XL [98], NF-RegNet-B1 [10]
@@ 6 Stem layer: 7$\times$7 conv w/ s2, ReLU, ResNet-50 [42], ResNet-101 [42], ResNeXt-50 [117],
8 Input resolution: 224, Stem layer: stack of 3$\times$3 convs, NFNet-L0222Customized models by [114]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [11]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [66]., ECA-NFNet-L0222Customized models by [114]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [11]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [66]., PoolFormer-M48 [121],
Group Conv, Final GAP, 2D SA ResNeSt-50 [126], ResNet-V2-50-D-EVOS [^2] Customized models by [114]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [11]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [66].,
ConvMixer-1536/20 [104]
9 ReLU, Input resolution: 224, DW Conv, BN, ViT-B (patch size 32) [30], R26+ViT-S [93], DLA-X-102 [119],
2D self-attention eSE-VoVNet-39 [63], ResNet-101-C [45], RegNetX-320 [81],
HRNet-W32 [110]
10 ReLU + Leaky ReLU, InRes: 256, CSPResNet-50 [109], CSPResNeXt-50 [109], CSPDarkNet-53 [8],
Stem layer: 7$\times$7 conv, CNN, Pooling at stem NF-ResNet-50 [10]
Figure 3: Importance of architectural components to network similarity. 13 components are sorted by the contribution to the similarities. The larger feature importance means the component contributes more to the network similarity.
Figure 4: Pairwise distances of spectral features. Rows and columns are sorted by the clustering index. More details are described in Sec. 0.D.3.
Clustering analysis.
@@ 10 ReLU + Leaky ReLU, InRes: 256, CSPResNet-50 [109], CSPResNeXt-50 [109], CSPDarkNet-53 [8],
We additionally provide a clustering analysis based on the architectural similarities. We construct a pairwise similarity graph with adjacency matrix $A$ between all 69 architectures where its vertex denotes an architecture, and its edge denotes the similarity between two networks. We perform the spectral clustering [75] on $A$ where the number of clusters $K$ is set to 10: We compute the Laplacian matrix of $A$, $L=D-A$ where $D$ is the diagonal matrix and its $i$-th component is $\sum_{j}A_{ij}$. Then, we perform K-means clustering on the $K$-largest eigenvectors of $L$. The pairwise distances of spectral features (i.e., 10-largest eigenvectors of $L$) of 69 neural architectures are shown in Fig. 4. The rows and columns of Fig. 4 are sorted by the clustering index (Tab. 1). More details with model names are described in Sec. 0.D.2. We can see the block-diagonal patterns, i.e., in-clusters similarities are more significant than between-clusters similarities. More details are in Sec. 0.D.3.
@@
Table 2: SAT within the same architecture. We compare the average similarity within the same architecture but trained with different procedures, “All” denotes the average similarity of 69 architectures.
Architecture ResNet-50 ViT-S
Init 4.23 4.21
Hparam 4.05 4.22
Tr. Reg. 3.27 3.44
@@
Table 3: Ensemble performance with diverse architectures. We report the error reduction rate by varying the number of ensembled models and the diversity of the ensemble models (related to Fig. 7(a)). “rand” indicates the random choice of models.
less diverse $\leftarrow$ # of clusters $\rightarrow$ more diverse
1 2 3 4 5 rand
# of models 2 7.13 7.84 7.78
3 10.17 10.84 11.20 11.11
4 11.70 12.45 12.80 13.00 12.90
5 12.58 13.41 13.79 13.99 14.11 14.01
4.2 The Relationship between Training Strategy and SAT
The architectural difference is not the only cause of the model diversity. We compare the impact by different architecture choices (e.g., ResNet and ViT) and by different training strategies while fixing the model architecture, as follows: Different initializations can affect the model training by the nature of the stochasticity of the training procedure. For example, Somepalli et al. [90] showed that the decision boundary of each architecture could vary by different initializations. We also consider different optimization hyper-parameters (e.g., learning rate, weight decay). Finally, we study the effect of different training regimes (e.g., augmentations, type of supervision). For example, the choice of data augmentation [125, 122] or label smoothing [96] can theoretically or empirically affect adversarial robustness [127, 77, 87, 21]. We also investigate the effect of supervision, such as self-supervision [15, 44, 37] or semi-weakly supervised learning [118]. Note that the training strategies inevitably contain the former ones. For example, when we train models with different training regimes, models have different initialization seeds and different optimization hyper-parameters. Comparing 69 different architectures also contains the effect of different initialization and optimization hyper-parameters and parts of different training regimes. This is necessary for achieving high classification performance.
@@ 4.2 The Relationship between Training Strategy and SAT
Tab. 2 shows the comparison of similarity scores between the same architecture but different learning methods (a smaller similarity means more diversity). We report two architectures, ResNet-50 and ViT-S, and their training settings are in Appendix 0.E. We also show the average SAT between all 69 architectures. In the table, we first observe that using different random initialization or different optimization hyper-parameters shows high correlations with each other (almost $\geq$ 4.2) while the average similarity score between various neural architectures is 2.73. In other words, the difference in initializations or optimization hyper-parameters does not significantly contribute to the model diversity.
Second, we observe that using different learning techniques remarkably affects SAT (3.27 for ResNet and 3.44 for ViT), but is not as significant as the architectural difference (2.73). Furthermore, the change of SAT caused by different initializations or hyper-parameters is less marked than the change caused by different architecture (Fig. 6). These observations provide two insights. First, the diversity resulting from various training strategies is not significant enough compared to the diversity of architecture. Second, designing new architecture is more efficient in achieving diverse models rather than re-training the same one.
Figure 5: Pairwise distance of spectral features by different optimizations. Init, LR, and WD are randomly chosen from models trained with different settings of initialization, learning rate, and weight decay in Tab. 2.
@@
Figure 6: Correlation between pairwise SAT and ensemble performance. The trend line and its 90% confidence interval are shown.
5 SAT Applications with Multiple Models
Here, we analyze how SAT is related to downstream tasks involving more than one model. First, we show that using more diverse models will lead to better ensemble performance. Second, we study the relationship between knowledge distillation and SAT. Furthermore, we can suggest a similarity-based guideline for choosing a teacher model when distilling to a specific architecture. Through these observations, we can provide insights into the necessity of diverse models.
5.1 Model Diversity and Ensemble
Settings.
@@ 5.1 Model Diversity and Ensemble
Results.
We first measure the 2-ensemble performances among the 69 architectures (i.e., the number of ensembles is $\binom{69}{2}=2346$). We plot the relationship between SAT and ERR in Fig. 6. We observe that there exists a strong negative correlation between the model similarity and the ensemble performance (Pearson correlation coefficient $-0.32$ with p-value $\approx 0$ and Spearman correlation $-0.32$ with p-value $\approx 0$, i.e., more diversity leads to better ensemble performance.
We also conduct $N$-ensemble experiments with $N\geq 2$ based on our clustering results in Tab. 1. We evaluate the average ERR of the ensemble of models from $k$ clusters, i.e., if $N=5$ and $k=3$, the ensembled models are only sampled from the selected 3 clusters while ignoring the other 7 clusters. We investigate the effect of model diversity and ensemble performance by examining $k=1\ldots N$ (i.e., larger $k$ denotes more diverse ensembled models). We report the result with ImageNet top-1 error and ERR in Fig. 7(a) and Fig. 7(b).
(a) Top-1 Error
(b) Error reduction rate
(c) “All wrong” samples ratio
Figure 7: Model diversity and ensemble performance. We report ensemble performances by varying the number of ensembled models ($N$) and the diversity of the models. The diversity is controlled by choosing the models from $k$ different clusters.
@@ Results.
Figure 8: Diversity by training techniques and ensemble. We report the the same metrics as Fig. 7 for various ResNet-50 and ViT-S models in Tab. 2.
In all metrics, we observe that the ensemble of more diverse models shows better performance. Interestingly, Fig. 7(b) shows that when the number of clusters for the model selection ($k$) is decreased, the ensemble performance by the number of ensembled models ($N$) quickly reaches saturation. Tab. 3 shows that the ensemble performances by choosing the most diverse models via SAT always outperform the random ensemble. Similarly, Fig. 7(c) shows that the number of wrong samples by all models is decreased by selecting more diverse models.
Training Strategy vs. Architecture in the ensemble scenario?
Remark that Tab. 2 showed that the different training strategies are not as effective as different architectures for diversity. To examine this on the ensemble scenario, we report the ensemble results of different training strategies, i.e., the same ResNet-50 and ViT-S in Tab. 2. For comparison with different architectures, we also report the ensemble of different architectures where all ensembled models are from different clusters (i.e., $N$=$k$ in Fig. 7). Fig. 8 shows that although using diverse training regimes (blue lines) improves ensemble performance compared to other techniques (red and green lines), the improvements by using different architectures (yellow lines) are more significant than the improvements by using different training regimes (blue lines) with large gaps.
(a) SAT vs. Flowers ensemble
(b) SAT vs. ImageNet ensemble
Figure 9: Cross-dataset SAT results. SAT measured on ImageNet also has a positive correlation with ensemble performances on Flowers-102 [76].
Figure 10: Empircial comparison. Relationship between the model similarity, including SAT, Somepalli et al. [90] and Tramèr et al. [103], and 2-ensemble performance.
Generalizability to other datasets.
We examine whether more diverse architectures in ImageNet SAT also lead to better ensemble performances on the other datasets. We fine-tuned all 69 architectures to the Flowers-102 dataset [76], and filter out low performing models (< 95% top-1 accuracy). After the filtering, we have 16 fine-tuned models. Using the fine-tuned models, we plot the relationship between the Flowers-102 ensemble performance and SAT score measured in ImageNet. Fig. 9 shows that SAT also highly correlates with Flowers ensemble performances, despite that SAT is measured on ImageNet. This experimental result supports that SAT similarity can be applied in a cross-domain manner.
Comparison of different similarity functions in the ensemble scenario.
Finally, we compare the impact of the choice of the similarity function and the ensemble performance when following our setting. We compare SAT with Somepalli et al. [90] and Tramèr et al. [103] on the 2-ensemble scenario with 8 out of 69 models due to the stability issue of Tramèr et al. [103]. Fig. 10 shows the relationship between various similarity functions and the 2-ensemble performance. We observe that SAT only shows a strong positive correlation (blue line), while the others show an almost random or slightly negative correlation. Finally, in Sec. 0.A.2, we compare SAT with Somepalli et al. [90] and a naive architecture-based clustering using our features under the same setting of Fig. 7 and 8. Similarly, SAT shows the best ensemble performance against the comparison methods.